OpenStudy (anonymous):

Factor 3x^2y^2 - xy^2 - 2y^2

OpenStudy (anonymous):

I belive that the answer is 3x - 2

ganeshie8 (ganeshie8):

First pull the GCF out

ganeshie8 (ganeshie8):

you have 3 terms, whats the common factor ? (if exists)

OpenStudy (anonymous):

its xy^2 right?

ganeshie8 (ganeshie8):

*last term have no x-terms

OpenStudy (anonymous):

Oh right so only y^2

ganeshie8 (ganeshie8):

Yep, pull that out

OpenStudy (anonymous):

3x^2 - x - 2

ganeshie8 (ganeshie8):

3x^2y^2 - xy^2 - 2y^2 y^2( 3x^2 - x - 2 )

OpenStudy (anonymous):

So now wouldn't that be the answer 3x^2 -x - 2

OpenStudy (anonymous):

I believe that you cant do anything else to it

ganeshie8 (ganeshie8):

you can still factor 3x^2 - x - 2

OpenStudy (anonymous):

Huh but by what?

ganeshie8 (ganeshie8):

factor by grouping

ganeshie8 (ganeshie8):

pick any two number such that : they multiply to -6 add up to -1

OpenStudy (anonymous):

umm -3 and -2?

OpenStudy (anonymous):

Or postive 3 and negative 2

ganeshie8 (ganeshie8):

+3 -2 is not -1

OpenStudy (anonymous):

What do you so with those numbers now?

ganeshie8 (ganeshie8):

the magic numbers are -3 and +2 : -3*2 = -6 -3 + 2 = -1

OpenStudy (anonymous):

what do i do with -3 and +2 when i have them

ganeshie8 (ganeshie8):

3x^2 - x - 2 split the second term as : -3x + 2x

ganeshie8 (ganeshie8):

3x^2 - x - 2 3x^2 - 3x + 2x - 2

ganeshie8 (ganeshie8):

factor the GCF from first two terms and last two terms separately

ganeshie8 (ganeshie8):

3x^2 - x - 2 3x^2 - 3x + 2x - 2 3x(x-1) + 2(x-1) (x-1)(3x+2)

ganeshie8 (ganeshie8):

the final factorization is : 3x^2y^2 - xy^2 - 2y^2 = y^2(x-1)(3x+2)

OpenStudy (anonymous):

Wouldn't that be (3x + 2)(x - 1)

ganeshie8 (ganeshie8):

yes, both are same

ganeshie8 (ganeshie8):

the final answer is ` y^2(x-1)(3x+2)` you must include `y^2` also, okay ?

OpenStudy (the_fizicx99):

Same thing, multiplication is commutative, order doesn't matter >.>

OpenStudy (anonymous):

Know what if you cant do the simplfying like in this umm x^2 + 6x + 9

OpenStudy (the_fizicx99):

Perfect square? (x + 3)^2 = (x)^2 + 2(x)(3) + (3)^2 Or, (x+3)(x+3) = x^2 + 3x + 3x + 9 = x^2 + 6x + 9

OpenStudy (anonymous):

Thank you guys :) Have a great day

OpenStudy (the_fizicx99):

Yw :>